OpenStudy (anonymous):

Write the equation of the quadratic function with roots 0 and 1 and a vertex at (1/2, -2/3)

OpenStudy (amistre64):

how do the roots helps us?

OpenStudy (anonymous):

I did the work and i start getting confused after -2/3 a (1/2)^2 - 1 (1/2 + 0) @amistre64

OpenStudy (amistre64):

roots tell us about factors x(x-1) will zero out when x=0, or x=1 and any multiple of this will also have the same zeros; since anything times 0 = 0 y = ax(x-1) now define a by using the given point x=1/2, y=-2/3 -2/3 = a 1/2 (1/2-1) -4/3 = a (-1/2) etc ...

OpenStudy (amistre64):

the quadratic of course might be of the form: y = ax^2 -ax, but im not sure how they want it presented.